OpenStudy (anonymous):

Abby wants to buy some fruit. She has $10.00 to spend. Apples (x) cost $0.50 a pound and bananas (y) cost $1.50 a pound. Which graph shows all of the ways she can spend up to $10.00 on apples and bananas? PLEASEEEEE HELPPPP! !! ASAP

OpenStudy (whpalmer4):

got graph? :-)

OpenStudy (anonymous):

YESSS

OpenStudy (whpalmer4):

We can write an equation for the ways she can spend up to $10 if apples are $0.50/lb and bananas are $1.50/lb. The price (in dollars) of \(x\) pounds of apples is \(0.50x\) The price (in dollars) of \(y\) pounds of bananas is \(1.50y\) Price of a mix of apples and bananas is just \(0.50x+1.50y\) We want to spend no more than $10, so \[0.50x+1.50y \le 10\] That's the inequality that shows the ways she can buy up to $10 worth of apples and bananas.

OpenStudy (whpalmer4):

Now you just need to figure out which graph corresponds to that inequality. The dividing line on the graph will be an equation, \(0.50x + 1.50y = 10\) and it will be a solid line, not dashed, because we have the inclusive inequality sign. Find the x and y intercepts of that line by putting y = 0 and x = 0 respectively and solving for the other. Plot those two points, draw a line through them, and you've got your graph mostly done. Now pick an easy point to test — (0,0) is my favorite — and see if it satisfies the inequality. If it does, shade the region containing the test point. If it doesn't, the region containing the test point is not the one to shade. Compare with the graphs in the answer choices and make your selection.

OpenStudy (anonymous):

OK THIS IS SUPER HARDD. !!! which side is it going to be shaded on ?

OpenStudy (whpalmer4):

I gave you a test point: (0,0) That means (x=0,y=0) Try that out in the inequality: \[0.50(x) + 1.50(y) \le 10\]\[0.50(0)+1.50(0)\le 10\]Is that true? If it is, shade the side of the line that has the point (0,0). If it isn't, shade the other side of the line.

OpenStudy (anonymous):

holy pellet in dont care about that i just want the answer

OpenStudy (anonymous):

holy pellet in dont care about that i just want the answer